# Your Token Becomes Worthless: Unveiling Rug Pull Schemes in Crypto Token via Code-and-Transaction Fusion Analysis

### v1

@@ 5.1. Experiment Setup
5.2. RQ1: Effectiveness of RPHunter
To evaluate the effectiveness of RPHunter, we compared its performance with five state-of-the-art Rug Pull detection methods. There are two selection criteria for these SOTA methods: (1) open-source availability or ease of replication, and (2) suitability for automated large-scale experiments. The five methods are Ma et al. (Ma et al., 2023), Lin et al. (Lin et al., 2024), Mazorra et al. (Mazorra et al., 2022) and Xia et al. (Xia et al., 2021) from academia and GoPlus (GoP, 2025) from industry. Ma et al. (Ma et al., 2023) proposed Pied-Piper, which integrates datalog analysis and fuzzing to detect some Rug Pulls. Lin et al. (Lin et al., 2024) proposed CRPWaner, the most recent academic tool that detects Contract-related Rug Pulls based on static analysis. We used these two methods as the baselines for rule-based methods which are open-sourced (CRP, 2025; pie, 2025). Mazorra et al. (Mazorra et al., 2022) and Xia et al. (Xia et al., 2021) utilized statistical transaction features to train scam token detection models. We used these two methods as the baseline for learning-based methods, replicated their models based on their paper, and conducted experiments on our ground-truth dataset. GoPlus (GoP, 2025) is a widely known commercial tool, providing APIs to detect scam tokens. We used GoPlus as the baseline for commercial methods.
Table 3. Performance comparison with SOTA methods
@@ 5.2. RQ1: Effectiveness of RPHunter
Pied-Piper (Ma et al., 2023) 58.8% 5.8% 1.7% 94.2% 10.5%
CRPWarner (Lin et al., 2024) 31.2% 56.2% 47.3% 43.8% 40.1%
Xia et al. (Xia et al., 2021) 70.8% 75.8% 12.1% 24.2% 73.2%
Mazorra et al. (Mazorra et al., 2022) 71.9% 79.4% 11.9% 20.6% 75.5%
GoPlus (GoP, 2025) 74.0% 78.2% 10.6% 21.4% 76.2%
@@
The results can be seen in Table 3, where RPHunter outperforms all baselines on the ground-truth dataset, achieving a precision of 95.3%, a recall of 93.8%, an FPR of 1.8%, an FNR of 6.2%, and an overall F1 score of 94.5%. However, Pied-Piper achieves the lowest FPR, since it only targets a small part of Rug Pull code risks. Due to limited detection capability, it naturally exhibits the lowest FPR at the cost of a extremely high FNR.
Impact of Code: In the dataset, there are 229 incidents that exhibit no obvious code risks and primarily initiate Rug Pulls through unlocked liquidity, which are common (Certik, 2025). This led to relatively poor performance by the static analysis tools, Pied-Piper and CRPWarner. In contrast, RPHunter not only captures semantic information from the code but also leverages early-stage transactions that reveal malicious behaviors. This combination enables effective detection of Rug Pulls, even for tokens without apparent code risks.
@@
Table 4. Impact of Transaction Volume on Performance
@@
5.3. RQ2: Time Overhead of RPHunter
Given a crypto token address, we evaluated the time overhead of RPHunter for detecting Rug Pull risks within the token. To accurately evaluate the time overhead, we tested each crypto token five times in our dataset and recorded the average time. As shown in Figure 6, the average time consumption results of each component of RPHunter is 20.69s for SRCG Construction, 6.24s for TFBG Construction, and 0.26s for Rug Pull Detection. Due to the parallel processing capabilities in constructing SRCG and TFBG, the average time overhead of RPHunter is 20.95s. This time overhead is acceptable for early Rug Pull detection, enabling proactive measures before financial losses occur. Most of the overhead of RPHunter comes from complex flow analysis.
Figure 5. Overhead of Each Component
Figure 6. Overhead Comparison with SOTA methods
We also calculate the time overhead for other SOTA methods on the same dataset. As shown in Figure 6, it is evident that rule-based methods (Ma et al., 2023; Lin et al., 2024) are time-consuming, taking 28.36s and 103.46s respectively for each crypto token. Alternatively, learning-based methods (Xia et al., 2021; Mazorra et al., 2022), which utilize machine learning, demonstrate a shorter detection time. The majority of its time overhead is spent on processing transaction data to extract statistical features. Learning-based methods are generally faster than rule-based methods, but lack code analysis. Furthermore, the commercial tool GoPlus (GoP, 2025) shows the shortest detection time in standard scenarios. However, when analyzing newly deployed tokens, GoPlus may take over 5 minutes to produce a detection result. In summary, RPHunter presents a balanced approach with an acceptable time overhead and offers a comprehensive analysis of code semantic information.
5.4. RQ3: Ablation Study
@@
Table 5. Ablation Study Results
@@
5.4.1. Effectiveness of extracting code risk information in SRCG
We conducted a variant experiment, denoted as $SRCG_{NoSR}$ (V2), which excludes the code risk information from the SRCG. As shown in Table 5, compared to the $SRCG$ (V1), the precision, recall, and F1 score of $SRCG_{NoSR}$ decrease by 6.1%, 7.4%, and 6.7%, respectively. The results highlight the $SRCG$ holds more fine-grained code risk semantic information in the graphical features and underscore the critical role of code risk information in improving detection performance within the SRCG.
5.4.2. Effectiveness of incorporating node and edge features in TFBG
We conducted two variant experiments by respectively removing node and edge features from the TFBG, denoted as $TFBG_{NoN}$ (V4) and $TFBG_{NoE}$ (V5). As shown in Table 5, compared with the $TFBG$ (V3) with node and edge features, the absence of node features in $TFBG_{NoN}$ leads to decreases in precision, recall, and F1 score by 5.3%, 17.6%, and 9.6% respectively. Similarly, the removal of edge features in $TFBG_{NoE}$ results in a drop of 22.0% in recall, and 11.2% in F1 score. The absence of node and edge features leads to a significant decrease in recall and an increase in the FNR, indicating that both node and edge features play an important role in capturing transaction behaviors. We did not conduct an experiment that removed both node and edge features, which strip all meaningful information of TFBG, rendering it ineffective for analysis.
5.4.3. Effectiveness of fusion analysis of SRCG and TFBG
@@
5.5. RQ4: Rug Pull in Real World
@@
To evaluate RPHunter’s performance on crypto tokens beyond our benchmark, we conducted preliminary experiments on the real-world Ethereum Mainnet. We collected and analyzed contracts deployed from May 2024 to July 2024, spanning blocks 19771560 to 20207949. Each contract’s bytecode was examined to determine if the public function signatures conform to the ERC20 token standards. Utilizing the methodologies described in Section 4, we constructed the respective SRCG and TFBG for each token and fed them into the trained model for Rug Pull detection.
@@
Ultimately, RPHunter successfully identified 4801 Rug Pull tokens. We sampled 247 open-source tokens based on a 99% confidence level and an 8% confidence interval (Con, 2025) for in-depth analysis to substantiate the detection results. This involved a thorough investigation of each contract’s transaction history and code patterns that are indicative of Rug Pull activities. Upon analysis, we found 23 false positives, and the overall true positive rate achieves 91%. Through a deeper investigation into false positive cases, we identified the main reason. In certain instances, scammers refrain from writing complex contracts or initiating additional transactions to attract investment, which is costly and cumbersome, increasing the risk of detection. Instead, they just leverage off-chain social media to attract investors and exit by removing liquidity. This liquidity issue is closely tied to off-chain trust mechanisms. We plan to explore this in the future. In addition, further analysis of these Rug Pull tokens reveals that 108 of these tokens existed for no more than 24 hours, indicating that Rug Pulls are executed with remarkable speed. The massive issuance of short-lived Rug Pull token not only highlights the prevalence of fraudulent schemes but also emphasizes the necessity of the detection framework. We discuss one case study found on real-world Ethereum Mainnet for illustration (MNH, 2025).
Figure 7. Case Study
@@
Case study on MNHA token. Figure 7 shows the Rug Pull procedure. As the token price increased, the scammers triggered the Address Restrict vulnerability (line 10). This tactic restricted the liquidity pool contract from processing transactions, effectively preventing users from swapping their MNHA tokens ($S_{1}$). Once the scammers permitted swap transaction (line 13, $S_{2}$), they quickly utilized unlocked liquidity to exchange WETH, draining the liquidity pool ($S_{3}$). The entire process from token deployment to the completion of the Rug Pull lasted only 8 minutes. RPHunter effectively identifies the Rug Pull before the liquidity extraction occurred.
5.6. Threats to Validity
Our study faces several threats to validity introduced by the methods. (1) We select a set of Rug Pull incidents as the ground-truth dataset. However, our benchmark may still include a certain sample bias. Further study on more Rug Pull incidents can help better evaluate RPHunter. (2) For code risk information extraction, the effectiveness of RPHunter is threatened by the insufficiently precise flow analysis (produced by Gigahorse (Grech et al., 2019)). Since our method only leverages Gigahorse to extract code risk information, this threat can be mitigated by integrating more advanced tools.
6. RELATED WORK
@@
Rug Pull Detection. Scam activities are pervasive in the blockchain ecosystem, posing serious threats to the credibility and security of decentralized systems. Various types of scams, such as Ponzi schemes, and phishing attacks, have been widely studied in recent years (Sun et al., 2020; Chen et al., 2021; Torres et al., 2019; Roy et al., 2024; Lin et al., 2021; Xia et al., 2020; Béres et al., 2021; Bartoletti et al., 2020; Liebau and Schueffel, 2019; Sureshbhai et al., 2020; Wu et al., 2020; Liang et al., 2024; Zhou et al., 2024; Lin et al., 2024). Recently, Rug Pull has emerged as one of the most severe and damaging forms of scams (Cernera et al., 2023; Huang et al., 2023; Sechting and Raschke, 2024). Some research (Xia et al., 2021; Mazorra et al., 2022) explored machine learning techniques to identify Rug Pull tokens, utilizing transactions and statistical features. Ma et al. (Ma et al., 2023) defined five distinct types of backdoors in smart contracts and employed a combination of datalog analysis and fuzzing techniques to detect these backdoors. Furthermore, Lin et al. (Lin et al., 2024) and Zhou et al. (Zhou et al., 2024) performed an empirical study on Rug Pull events and exposed Rug Pull risks utilizing domain-specific datalog analysis. Commercial tools like GoPlus (GoP, 2025), TokenSniffer (Tok, 2025) audit token contracts based on the source code or EVM bytecode of smart contracts. In contrast to existing methods that focus on either code patterns or transaction behaviors. RPHunter integrates code and transaction information for comprehensive Rug Pull detection.
@@
Neural Networks for Vulnerability Detection. Various tools have been proposed to detect smart contract vulnerabilities using deep neural networks (Wu et al., 2020; Wang et al., 2020; Wu et al., 2021; Hu et al., 2021; Gopali et al., 2022; Zhuang et al., 2021; Cheng et al., 2024; Jeon et al., 2021; Wang et al., 2024; Sendner et al., 2023). Jeon et al. (Jeon et al., 2021) extracted code fragments from smart contracts and detected vulnerable code patterns using a pre-trained BERT model. Nevertheless, these sequence-based methods assume that the smart contract is a sequence of tokens without considering the graph structure of the contracts (Huang et al., 2021; Zeng et al., 2022). Graph neural networks have been proven to excel in vulnerability detection by leveraging the graph structure of code, which represents both the control flow and data dependency among different instructions (Liu et al., 2021; Cheng et al., 2024). Zhuang et al. (Zhuang et al., 2021) and Liu et al. (Liu et al., 2021) constructed code semantic graphs from source code and proposed TMP (Zhuang et al., 2021) and AME (Liu et al., 2021) to detect vulnerabilities, respectively. SCVHunter (Luo et al., 2024) designed a heterogeneous semantic graph based on intermediate representations and proposed a heterogeneous graph attention network for detection. PonziGuard (Liang et al., 2024) collected the contract’s runtime behavior using fuzzing techniques and established the contract runtime behavior graph to accurately detect Ponzi schemes.
7. CONCLUSION
@@
In this paper, we propose RPHunter, a novel Rug Pull detection framework that integrates code and transaction information, modeling these information within SRCG and TFBG for feature learning to unveil Rug Pull schemes. We evaluate RPHunter on a manually-labeled dataset of 645 Rug Pull tokens and an open-source dataset of 1675 benign tokens. The results show that RPHunter effectively detects Rug Pull tokens on the dataset, outperforming five state-of-the-art methods with an acceptable time overhead. Moreover, we apply RPHunter on the Ethereum Mainnet. It has identified 4801 Rug Pull tokens achieving a precision of 91%, which demonstrates RPHunter’s effectiveness in real-world scenarios.

### v2

@@ V-A Experiment Setup
V-B RQ1: Effectiveness of RPHunter
To evaluate the effectiveness of RPHunter, we compared its performance with five state-of-the-art Rug Pull detection methods. There are two selection criteria for these SOTA methods: (1) open-source availability or ease of replication, and (2) suitability for automated large-scale experiments. The five methods are Ma et al. [3], Lin et al. [4], Mazorra et al. [7] and Xia et al. [6] from academia and GoPlus [32] from industry. Ma et al. [3] proposed Pied-Piper, which integrates datalog analysis and fuzzing to detect some Rug Pulls. Lin et al. [4] proposed CRPWaner, the most recent academic tool that detects Contract-related Rug Pulls based on static analysis. We used these two methods as the baselines for rule-based methods which are open-sourced [33, 34]. Mazorra et al. [7] and Xia et al. [6] utilized statistical transaction features to train scam token detection models. We used these two methods as the baseline for learning-based methods, replicated their models based on their paper, and conducted experiments on our ground-truth dataset. GoPlus [32] is a widely known commercial tool, providing APIs to detect scam tokens. We used GoPlus as the baseline for commercial methods.
TABLE III: Performance comparison with SOTA methods
@@ V-B RQ1: Effectiveness of RPHunter
Pied-Piper [3] 58.8% 5.8% 1.7% 94.2% 10.5%
CRPWarner [4] 31.2% 56.2% 47.3% 43.8% 40.1%
Xia et al. [6] 70.8% 75.8% 12.1% 24.2% 73.2%
Mazorra et al. [7] 71.9% 79.4% 11.9% 20.6% 75.5%
GoPlus [32] 74.0% 78.2% 10.6% 21.4% 76.2%
@@
The results can be seen in Table III, where RPHunter outperforms all baselines on the ground-truth dataset, achieving a precision of 95.3%, a recall of 93.8%, an FPR of 1.8%, an FNR of 6.2%, and an overall F1 score of 94.5%. However, Pied-Piper achieves the lowest FPR, since it only targets a small part of Rug Pull code risks. Due to limited detection capability, it naturally exhibits the lowest FPR at the cost of a extremely high FNR.
Impact of Code: In the dataset, there are 229 incidents that exhibit no obvious code risks and primarily initiate Rug Pulls through unlocked liquidity, which are common [35]. This led to relatively poor performance by the static analysis tools, Pied-Piper and CRPWarner. In contrast, RPHunter not only captures semantic information from the code but also leverages early-stage transactions that reveal malicious behaviors. This combination enables effective detection of Rug Pulls, even for tokens without apparent code risks.
@@
TABLE IV: Impact of Transaction Volume on Performance
@@
V-C RQ2: Time Overhead of RPHunter
Given a crypto token address, we evaluated the time overhead of RPHunter for detecting Rug Pull risks within the token. To accurately evaluate the time overhead, we tested each crypto token five times in our dataset and recorded the average time. As shown in Fig. 6, the average time consumption results of each component of RPHunter is 20.69s for SRCG Construction, 6.24s for TFBG Construction, and 0.26s for Rug Pull Detection. Due to the parallel processing capabilities in constructing SRCG and TFBG, the average time overhead of RPHunter is 20.95s. This time overhead is acceptable for early Rug Pull detection, enabling proactive measures before financial losses occur. Most of the overhead of RPHunter comes from complex flow analysis.
Figure 5: Overhead of Each Component
Figure 6: Overhead Comparison with SOTA methods
We also calculate the time overhead for other SOTA methods on the same dataset. As shown in Fig. 6, it is evident that rule-based methods [3, 4] are time-consuming, taking 28.36s and 103.46s respectively for each crypto token. Alternatively, learning-based methods [6, 7], which utilize machine learning, demonstrate a shorter detection time. The majority of its time overhead is spent on processing transaction data to extract statistical features. Learning-based methods are generally faster than rule-based methods, but lack code analysis. Furthermore, the commercial tool GoPlus [32] shows the shortest detection time in standard scenarios. However, when analyzing newly deployed tokens, GoPlus may take over 5 minutes to produce a detection result. In summary, RPHunter presents a balanced approach with an acceptable time overhead and offers a comprehensive analysis of code semantic information.
V-D RQ3: Ablation Study
@@
TABLE V: Ablation Study Results
@@
V-D1 Effectiveness of extracting code risk information in SRCG
We conducted a variant experiment, denoted as $SRCG_{NoSR}$ (V2), which excludes the code risk information from the SRCG. As shown in Table V, compared to the $SRCG$ (V1), the precision, recall, and F1 score of $SRCG_{NoSR}$ decrease by 6.1%, 7.4%, and 6.7%, respectively. The results highlight the $SRCG$ holds more fine-grained code risk semantic information in the graphical features and underscore the critical role of code risk information in improving detection performance within the SRCG. Notably, the $SRCG$ (v1) achieves a high recall of 89.1%, even though our earlier empirical study identifies 229 Rug Pulls that exhibit no explicit code risks. Our analysis suggests that while these contracts may lack clearly malicious code, they often exhibit subtle structural differences from normal contracts. The SRCG captures these fine-grained patterns, and the Bert-based encoder effectively learns these deviations during training. As a result, the high recall of V1 stems not only from detecting known risky patterns, but also from recognizing latent structural and semantic deviations that associated with Rug Pull contracts.
V-D2 Effectiveness of incorporating node and edge features in TFBG
We conducted two variant experiments by respectively removing node and edge features from the TFBG, denoted as $TFBG_{NoN}$ (V4) and $TFBG_{NoE}$ (V5). As shown in Table V, compared with the $TFBG$ (V3) with node and edge features, the absence of node features in $TFBG_{NoN}$ leads to decreases in precision, recall, and F1 score by 5.3%, 17.6%, and 9.6% respectively. Similarly, the removal of edge features in $TFBG_{NoE}$ results in a drop of 22.0% in recall, and 11.2% in F1 score. The absence of node and edge features leads to a significant decrease in recall and an increase in the FNR, indicating that both node and edge features play an important role in capturing transaction behaviors. We did not conduct an experiment that removed both node and edge features, which strip all meaningful information of TFBG, rendering it ineffective for analysis.
V-D3 Effectiveness of fusion analysis of SRCG and TFBG
@@
V-E RQ4: Rug Pull in Real World
@@
To evaluate RPHunter’s performance on crypto tokens beyond our benchmark, we conducted preliminary experiments on the real-world Ethereum Mainnet. We collected and analyzed contracts deployed from May 1 2024 to May 29 2024, spanning blocks 19771560 to 19972630. Each contract’s bytecode was examined to determine if the public function signatures conform to the ERC20 token standards. Utilizing the methodologies described in Section IV, we constructed the respective SRCG and TFBG for each token and fed them into the trained model for Rug Pull detection.
Ultimately, RPHunter successfully identified 4801 Rug Pull tokens from 9528 ERC20 Tokens. This proportion is alarmingly high, yet consistent with prior industry findings, for example, blockchain security firm Certik [35] has reported that one in every two new tokens on Ethereum is involved in scam. We sampled 247 open-source tokens based on a 99% confidence level and an 8% confidence interval [36] for in-depth analysis to substantiate the detection results. This involved a thorough investigation of each contract’s code patterns and transaction behaviors. Specifically, we first retrieve the source code from blockchain explorer and perform double-check manual auditing to identify the presence of predefined code risks. If any such risks were confirmed, we flagged the token as a potential Rug Pull. Then we further analyzed each token’s price trajectory using CoinMarketCup [1] and associated transaction records. We investigated whether the token experienced a sudden and irreversible price crash, and if such a crash was tied to suspicious liquidity withdrawals or concentrated token transfers, thereby corroborating the presence of a Rug Pull event.
@@
Upon analysis, we found 23 false positives, and the overall true positive rate achieves 90.7%. Through a deeper investigation into false positive cases, we identified the main reason. In certain instances, scammers refrain from writing complex contracts or initiating additional transactions to attract investment, which is costly and cumbersome, increasing the risk of detection. Instead, they just leverage off-chain social media to attract investors and exit by removing liquidity. This liquidity issue is closely tied to off-chain trust mechanisms. We plan to explore this in the future. In addition, further analysis of these Rug Pull tokens reveals that 108 of these tokens existed for no more than 24 hours, indicating that Rug Pulls are executed with remarkable speed. The massive issuance of short-lived Rug Pull token not only highlights the prevalence of fraudulent schemes but also emphasizes the necessity of the detection framework. We discuss one case study found on real-world Ethereum Mainnet for illustration [37].
Figure 7: Case Study
@@
Case study on MNHA token. Fig. 7 shows the Rug Pull procedure. As the token price increased, the scammers triggered the Address Restrict vulnerability (line 10). This tactic restricted the liquidity pool contract from processing transactions, effectively preventing users from swapping their MNHA tokens ($S_{1}$). Once the scammers permitted swap transaction (line 13, $S_{2}$), they quickly utilized unlocked liquidity to exchange WETH, draining the liquidity pool ($S_{3}$). The entire process from token deployment to the completion of the Rug Pull lasted only 8 minutes. RPHunter effectively identifies the Rug Pull before the liquidity extraction occurred.
V-F Threats to Validity
Internal Threats. The effectiveness of RPHunter is threatened by the insufficiently precise flow analysis (produced by Gigahorse [26]). Since our method only leverages Gigahorse to extract code risk information, this threat can be mitigated by integrating more advanced tools. Additionally, during the real-world validation on the Ethereum Mainnet, we rely on manual inspection to determine whether a contract constitutes a Rug Pull. This subjective process may introduce labeling inconsistencies or human error. To mitigate this threat, we validated token data through trusted platform CoinMarketCap and performed a double-check mechanism to ensure accuracy.
External Threats. Our ground-truth dataset is constructed by manually collecting and analyzing a set of Rug Pull incidents. Despite our efforts to gather a comprehensive set of known cases, the process may still introduce sample bias or contain inaccurate labeling. This could affect the effectiveness of RPHunter. To mitigate this threat, we sourced incidents from multiple channels and recruited a panel of three domain experts to review and verify each case, thereby increasing the reliability of datasets.
VI RELATED WORK
@@
Rug Pull Detection. Scam activities are pervasive in the blockchain ecosystem, posing serious threats to the credibility and security of decentralized systems. Various types of scams, such as Ponzi schemes, and phishing attacks, have been widely studied in recent years [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 5, 4]. Recently, Rug Pull has emerged as one of the most severe and damaging forms of scams [50, 51, 52]. Some research [6, 7] explored machine learning techniques to identify Rug Pull tokens, utilizing transactions and statistical features. Ma et al. [3] defined five distinct types of backdoors in smart contracts and employed a combination of datalog analysis and fuzzing techniques to detect these backdoors. Furthermore, Lin et al. [4] and Zhou et al. [5] performed an empirical study on Rug Pull events and exposed Rug Pull risks utilizing domain-specific datalog analysis. Commercial tools like GoPlus [32], TokenSniffer [53] audit token contracts based on the source code or EVM bytecode of smart contracts. In contrast to existing methods that focus on either code patterns or transaction behaviors. RPHunter integrates code and transaction information for comprehensive Rug Pull detection.
@@
Neural Networks for Vulnerability Detection. Various tools have been proposed to detect smart contract vulnerabilities using deep neural networks [54, 55, 56, 57, 58, 59, 60, 61, 62, 63]. Jeon et al. [61] extracted code fragments from smart contracts and detected vulnerable code patterns using a pre-trained BERT model. Nevertheless, these sequence-based methods assume that the smart contract is a sequence of tokens without considering the graph structure of the contracts [64, 65]. Graph neural networks have been proven to excel in vulnerability detection by leveraging the graph structure of code, which represents both the control flow and data dependency among different instructions [66, 60]. Zhuang et al. [59] and Liu et al. [66] constructed code semantic graphs from source code and proposed TMP [59] and AME [66] to detect vulnerabilities, respectively. SCVHunter [67] designed a heterogeneous semantic graph based on intermediate representations and proposed a heterogeneous graph attention network for detection. PonziGuard [49] collected the contract’s runtime behavior using fuzzing techniques and established the contract runtime behavior graph to accurately detect Ponzi schemes.
VII CONCLUSION
@@
In this paper, we propose RPHunter, a novel Rug Pull detection framework that integrates code and transaction information, modeling these information within SRCG and TFBG for feature learning to unveil Rug Pull schemes. We evaluate RPHunter on a manually-labeled dataset of 645 Rug Pull tokens and an open-source dataset of 1675 benign tokens. The results show that RPHunter effectively detects Rug Pull tokens on the dataset, outperforming five state-of-the-art methods with an acceptable time overhead. Moreover, we apply RPHunter on the Ethereum Mainnet. It has identified 4801 Rug Pull tokens achieving a precision of 90.7%, which demonstrates RPHunter’s effectiveness in real-world scenarios.